OpenStudy (anonymous):

How many different five-digit numbers can be formed using three 2s and two 5s?

OpenStudy (amistre64):

theres a counting rule for this:\[\frac{(a+b+c+...+k)!}{a!b!c!...k!}\]

OpenStudy (anonymous):

never heard of this rule, can you elaborate please?

OpenStudy (amistre64):

that was the elaboration

OpenStudy (amistre64):

we have 2+3 elements that we are trying to order we have 2 of one element, and 3 of the other

OpenStudy (amistre64):

how many ways are there to arrange: aaabbbcc? 8 elements, 3 are the same and 3 are the same and 2 are the same \[\frac{(3+3+2)!}{3!3!2!}\]

OpenStudy (anonymous):

oh okay i get it, one sec let me try it with question

OpenStudy (anonymous):

Got it thanks :D

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

oh btw what is this rule called?

OpenStudy (anonymous):

oh wait nvm i found the rule in my book, in the book is actually \[n!/a!b!c!...\]

OpenStudy (amistre64):

hmm, ive always known it as the counting rule .... http://www.hamilton.ie/ollie/EE304/Counting.pdf around slide 11

OpenStudy (amistre64):

yeah, its in a book too :)

OpenStudy (amistre64):

the n part is the number of elements, which is just the sumof its parts

OpenStudy (anonymous):

Right, anyways thanks for the help :)

OpenStudy (amistre64):

yw

OpenStudy (amistre64):

its a more generalized version of the permuation rule that we start with. take the the example: abcde how many ways can we arrange these? 5! ways of course but it fits in the formula as:\[\frac{(1+1+1+1+1)!}{1!1!1!1!1!}=\frac{5!}{1}\]